OpenStudy (anonymous):

Describe the function's level curves: f(x,y)=ln(x^2+y^2)

OpenStudy (experimentx):

i don't see a curve ... it's a fucntion of x and y, if f=z, you get a surface

OpenStudy (anonymous):

The set of points in the plane where a function f(x,y) has a constant value f(x,y)=c is called a level curve of f. So yes there are level curves

OpenStudy (experimentx):

then i guess so, however then it's a equation of circle

OpenStudy (anonymous):

how so, doesent the ln function contribute anythings

OpenStudy (experimentx):

with center 0,0 and radius C if the constant is lnC

OpenStudy (anonymous):

man, im just not geting it. How is it a circle

OpenStudy (experimentx):

f(x,y)=ln(x^2+y^2) = K = some ln(C) or x^2+y^2 = C which is circle, oops radius sqrt(c)